OpenStudy (anonymous):

A terrorist shoots at a police. The probability of bullet not hitting the police is 0.4. The probability of police dying after hitting the bullet is 0.6. What is the probability of the polic not dying after a shot?

OpenStudy (anonymous):

"The probability of police dying after hitting the bullet is 0.6." so if terrorist=1, then police life=0.4 "The probability of bullet not hitting the police is 0.4" so terrorist=0.6 0.6 x 0.4 = 0.24

OpenStudy (anonymous):

24% chance of survival if terrorist shoots at police

OpenStudy (anonymous):

no 0.64

OpenStudy (anonymous):

see before hit 0.4 -after hit 0.6=0.2

OpenStudy (anonymous):

0.6-0.4=0.2 not hit=0.4+0.2=0.6 i got that ans but answer is 0.64

OpenStudy (anonymous):

the probability of police surviving is a subset of being hit that's why I multiplied them

OpenStudy (anonymous):

ahhhh

OpenStudy (anonymous):

you're right: when the terrorist doesn't even hit them, that's another way to survive

OpenStudy (anonymous):

they can survive after being hit (0.24) or they can survive by not being hit (0.4)

OpenStudy (anonymous):

0.24+0.4=0.64 right

OpenStudy (anonymous):

yes, my first post calculates the chance of the police surviving with the 60% chance of being hit of course you have to give police another relief if the 40% of terrorist missing occur